# RT-H: Action Hierarchies Using Language

### v1

@@ II Related Work
Language has also been used to create hierarchy in multi-task learning. When tackling long horizon instructions, many recent approaches use LLMs or VLMs to decompose long horizon instructions into a sequence of tasks specified in language [1, 31, 32, 33, 16, 34]. Usually, scripted or individually trained policies are used to execute these tasks, limiting scalability. To learn long horizon policies end-to-end, Hu and Clune train a model to first predict language tasks and then predict actions conditioned on those tasks [35]. This approach is similar to RT-H but exists one level higher in the action hierarchy: they do not label or learn from fine-grained language motions. A few works explore the usage of more fine-grained language. Sharma et al. use an LLM to decompose tasks into a sequence of motion primitives [36]. Yet, these motion primitives are hard-coded and lack the contextuality that is required in more complex settings. RT-H learns language motions in the context of both the task and the scene, enabling better policies and more contextual corrections.
@@ V Experiments
Figure 3: Results on Diverse+Kitchen multi-task dataset, consisting of eight challenging evaluation tasks. RT-H outperforms RT-2 by 15% on average, getting higher performance on 6/8 of the tasks. Replacing language with class labels (RT-H-OneHot) drops performance significantly. Using action clusters via K-Means [53] instead of the automated motion labeling procedure leads to a minor drop in performance as well (RT-H-Cluster), demonstrating the utility of language motions as the intermediate action layer.
@@ V-A RT-H on Diverse Multi-Task Datasets
Train Dataset
Eval Dataset
RT-2 RT-H-Joint RT-H RT-H (GT)
Kitchen
Kitchen
30.2 28.22 24.9 17.9
D+K
Diverse
27.7 25.44 23.6 17.8
@@ V-B Contextual & Flexible Language Motions
Figure 6: Results for Corrections on models trained on the Diverse+Kitchen multi-task dataset, for the same eight evaluation tasks as in Fig. 3. RT-2-IWR is trained on teleoperation corrections from rolling out RT-2, while RT-H-Intervene is trained on language motion corrections from rolling out RT-H. RT-H-InterveneAction is trained on both language motion and action correction data. We see RT-H-Intervene both improves upon RT-H and substantially outperforms RT-2-IWR, suggesting that language motions are a much more sample efficient space to learn corrections than teleoperated actions. RT-H-InterveneAction performs better than RT-H, but fine-tuning actions sometimes leads to policy degeneration, since actions produced by RT-H during intervention can be suboptimal.
@@ V-D Generalization
pick
move
Average
RT-2
60
50
55
RT-H
70
60
65

### v2

@@ II Related Work
Language has also been used to create hierarchy in multi-task learning. When tackling long horizon instructions, many recent approaches use LLMs or VLMs to decompose long horizon instructions into a sequence of tasks specified in language [1, 31, 32, 33, 16, 34]. Usually, scripted or individually trained policies are used to execute these tasks, limiting scalability. To learn long horizon policies end-to-end, Hu and Clune train a model to first predict language tasks and then predict actions conditioned on those tasks [35]. This approach is similar to RT-H but exists one level higher in the action hierarchy: they do not label or learn from fine-grained language motions. A few works explore the usage of more fine-grained language, for example predefined motion primitives or by predicting object dynamics from verbs [36, 37]. Sharma et al. use an LLM to decompose tasks into a sequence of motion primitives [36]. Yet, these motion primitives are hard-coded and lack the contextuality that is required in more complex settings. RT-H learns language motions in the context of both the task and the scene, enabling better policies and more contextual corrections.
@@ V Experiments
Figure 3: Results on Diverse+Kitchen multi-task dataset, consisting of eight challenging evaluation tasks. 95% Wilson Score confidence intervals [54] are shown on the average success rates (left). RT-H outperforms RT-2 by 15% on average, getting higher performance on 6/8 of the tasks. Replacing language with class labels (RT-H-OneHot) drops performance significantly. Using action clusters via K-Means [55] instead of the automated motion labeling procedure leads to a minor drop in performance as well (RT-H-Cluster), demonstrating the utility of language motions as the intermediate action layer.
@@ V-A RT-H on Diverse Multi-Task Datasets
Train Dataset Eval Dataset RT-2 RT-H-Joint RT-H RT-H (GT)
Kitchen Kitchen 30.2 28.22 24.9 17.9
D+K Diverse 27.7 25.44 23.6 17.8
@@ V-B Contextual & Flexible Language Motions
Figure 6: Results for Corrections on models trained on the Diverse+Kitchen multi-task dataset, for the same eight evaluation tasks as in Fig. 3. 95% Wilson Score confidence intervals [54] are shown on the average success rates (left). RT-2-IWR is trained on teleoperation corrections from rolling out RT-2, while RT-H-Intervene is trained on language motion corrections from rolling out RT-H. RT-H-InterveneAction is trained on both language motion and action correction data. We see RT-H-Intervene both improves upon RT-H and substantially outperforms RT-2-IWR, suggesting that language motions are a much more sample efficient space to learn corrections than teleoperated actions. RT-H-InterveneAction performs better than RT-H, but fine-tuning actions sometimes leads to policy degeneration, since actions produced by RT-H during intervention can be suboptimal.
@@ V-C Training on Online Corrections
Failure Modes: RT-H demonstrates performance boosts on a wide variety of tasks, however the action hierarchy paradigm does lead to interesting failure modes. First, systemic language motion prediction errors can often confuse the action prediction, leading to oscillatory or incorrect behaviors where a flat model might not exhibit similar behaviors. For example, we occasionally noticed the robot would get stuck after closing the gripper by continuing to predict “close the gripper”, likely due to subtle observation changes between “close the gripper” frames and future frames. Thankfully, these issues are easy to debug and resolve, either through adjusting the automated labeling procedure or through intervention as we have shown.
Second, when collecting language motion corrections, the human operator is limited by the performance of the underlying action prediction model. When language motion corrections do not result in the correct behavior, it might become frustrating for the operator. For example if the operator asks the robot to “move arm left” over the bowl in the middle left row of Fig. 4, but the robot overshoots the bowl, this can make the process slower than teleoperation. This failure mode rarely happens for in-distribution tasks, but as tasks diverge from the data distribution, it becomes more likely.
@@ V-D Generalization
pick move Average
RT-2 60 50 55
RT-H 70 60 65
@@ VI Conclusion
Acknowledgments
We would like to thank Grecia Salazar, Deeksha Manjunath, Clayton Tan, Yansong Pang, Jornell Quimbao, Tran Pham, Utsav Malla, April Zitkovich, and Elio Prado for their help with dataset collection and on robot evaluation. We would also like to thank the greater Google DeepMind team for their feedback and contributions.